OpenStudy (highschoolmom2010):

Let A and B be independent events, C and D be mutually exclusive events, and F and G be overlapping events. Calculate the following probabilities if P(A)=25%, P(B)=25%, P(C)=10%, P(D)=40%, P(F)=80%, P(G)=50%, and P(F and G)=40%. find P(not A)

OpenStudy (anonymous):

P(not A) = 1 - P(A) = 1 - 0.25 = 0.75

OpenStudy (anonymous):

or 75%

OpenStudy (highschoolmom2010):

oh wow i think i way over thought that problem thanks @BangkokGarrett

OpenStudy (anonymous):

Yeah...are there more questions involving B C D F and G ?

OpenStudy (highschoolmom2010):

yes

OpenStudy (highschoolmom2010):

P (C or D) P ((not F) nd (not G))

OpenStudy (highschoolmom2010):

sorry supposed to say and*

OpenStudy (anonymous):

Since they're mutually exclusive you can just add them together P (C or D) = P(C) + P(D) = 0.1 + 0.4 = 0.5 or 50%

OpenStudy (highschoolmom2010):

i got that one i am kinda confused on the second one

OpenStudy (anonymous):

The second one is 1 - [P(F) + P(G) - P(F and G)] 1 - [0.8 + 0.5 - 0.4] 1 - 0.9 0.1

OpenStudy (highschoolmom2010):

the second one says NOT f and NOT g ........

OpenStudy (anonymous):

yes... i know... look at a Venn diagram We want the area outside those circles The area inside is 0.9 if you add the pieces The area outside is then just 1 - 0.9